# Emergence of non-ergodic multifractal quantum states in geometrical fractals

### v3

@@ Results
Left panel: Singularity spectrum $f(\alpha)$ for the representative states: extended (GS, red), localized (LOC, blue), and multifractal (NEM, orange). The spectrum for the Gaussian Orthogonal Ensemble (GOE, black line) is displayed as a reference for ideal extended states. Differences in the peak position and the width of the $f(\alpha)$ curve distinguish the localization classes. Right panel: Finite-size scaling of $f(\alpha)$ for the NEM case study. Data comparing different generation ranges show great agreement, supporting its multifractal nature.
@@
Through the preceding finite-size scaling analysis of these exemplary cases, one can examine the localization properties across different groups of eigenstates, potentially unveiling a variety of degrees of multifractality. The SM contains data for other NEMs, displaying different types of multifractal behavior, evident in their $f(\alpha)$ spectra—for example, in the width or the shift of the peak. Some of these NEMs can retain metallic or insulating characteristics [53, 54] and may also exhibit spatial structures that resemble bulk and edge modes observed in artificial fractal systems [34, 35].
@@
It is important to emphasize that, unlike systems described by Random Matrix Theory where an ensemble average over disorder realizations is usually performed, our system is deterministic and inherently disorder-free. Thus, $\mathcal{F}(\omega,E)$ directly measures the intrinsic correlations for our specific Hamiltonian between the densities of eigenstates $\psi_{i}$ and $\psi_{j}$ (with eigenvalues $\lambda_{i}$ and $\lambda_{j}$, respectively) that are separated by an energy $\omega$, with their average energy $(\lambda_{i}+\lambda_{j})/2$ centered around $E$. This approach provides an experimentally accessible probe to reveal and discriminate between extended, localized, and NEM signature states, particularly through its characteristic scaling behavior at criticality.
@@
Figure 6 illustrates the behavior of $\mathcal{F}(\omega,E)$ for the previously selected representative extended, localized, and NEM states, computed for the $6$-th generation SG. As predicted by theory, the two-point correlator for the extended state (red) and the localized state (blue) exhibits a plateau across several orders of magnitude at small energy separation in $\omega$. The intensity of this plateau reflects the degree of eigenstate overlap: it is relatively low and broad for extended states, due to their diffuse nature across the system, and significantly higher for localized states, which are concentrated in confined regions. At large energy separations in $\omega$, the correlator for both these types of states approaches an uncorrelated regime. Specifically, for localized states the decay is exponential. In this large-$\omega$ limit, $\mathcal{F}(\omega,E)$ tends towards unity (i.e., $\mathcal{F}\approx 1$), a value consistent with representative GOE statistics.
@@
Figure 6: Log-log plot of the two-point density-density correlator $\mathcal{F}(\omega,E)$ versus energy separation $\omega$. Data are shown for three characteristic states from a $6$-th generation Sierpiński Gasket: the extended ground state (GS, $E\simeq-3.667$, red markers), a localized state (LOC, $E\simeq-2.184$, blue markers), and a non-ergodic multifractal state (NEM, $E\simeq 1.509$, orange markers). The dashed black line indicates the Gaussian Orthogonal Ensemble (GOE) reference for uncorrelated systems, corresponding to $\mathcal{F}(\omega,E)\approx 1$. While both extended and localized states exhibit broad plateaus at small $\omega$ indicating non-vanishing correlations, their respective plateau heights are markedly different. In contrast, the NEM state features a narrower plateau (expected to become singular point in the thermodynamic limit) followed by a characteristic power-law decay, $\mathcal{F}(\omega,E)\propto\omega^{-\mu}$, a fingerprint of multifractality. Lines are guide to eye.
@@
The two-point correlator for the NEM state (orange) displays a distinctly shorter plateau, followed by a characteristic power-law decay with energy separation $\omega$, described by $\mathcal{F}(\omega,E)\propto\omega^{-\mu}$, before approaching the uncorrelated regime. Such power-law scaling, persisting across a significant range of $\omega$, is a hallmark of multifractal states and reflects their inherent scale-invariance [53, 56].
@@
To quantitatively verify this multifractal scaling and determine the correlation dimension $D_{2}$, we analyze the system-size dependence of $\mathcal{F}(\omega,E;N)$. Left panel of Figure 7 presents data for the NEM state at $E\simeq 1.509$ across multiple system sizes. Data for $g=5-7$ were obtained using the full spectrum (circles, solid lines), while for $g=8-9$, the correlator was evaluated using the available eigenstates computed via subspace diagonalization methods (triangles, dashed lines). As the system size $N$ increases, the $\mathcal{F}(\omega,E;N)$ curves exhibit improved convergence in their power-law decay region.
@@
Based on Chalker’s scaling hypothesis for critical systems [53], the exponent $\mu(N)$ is related to a size-dependent correlation dimension $D_{2}(N)$ by $\mu(N)\approx 1-D_{2}(N)$. Thus, the value for $D_{2}$ in the thermodynamic limit is obtained by extrapolating these finite-size $D_{2}(N)$ values against $1/\ln N$. The finite-size scaling data are plotted in the right panel of Fig. 7 and yield the fractal dimension: $D_{2}=0.655\pm 0.02$. This value is in excellent agreement, within the stated uncertainties, with the value $D_{2}\simeq 0.652$ obtained independently from IPR-based eigenstate statistics. This consistency between two distinct methods strongly corroborates the emergence of multifractal states.
@@
Left panel: Log-log plot of $\mathcal{F}(\omega,E;N)$ versus energy separation $\omega$ for Sierpiński Gasket generations $g=5-7$ (circles, solid lines, via exact diagonalization) and $g=8-9$ (triangles, dashed lines, via subspace diagonalization). Curves show convergence in the power-law decay regime $\mathcal{F}\propto\omega^{-\mu(N)}$. The dashed black line represents the GOE uncorrelated limit ($\mathcal{F}=1$).
Right panel: Finite-size scaling extrapolation of the correlation dimension $D_{2}(N)\approx 1-\mu(N)$ against $1/\ln(N)$. Blue circles are the $D_{2}(N)$ values derived from the decay exponents $\mu(N)$ for each generation. The dash-dot line is a linear fit, with its $y$-intercept (red circle) collapsing to thermodynamic limit value $D_{2}=0.655\pm 0.022$. Statistical errors are included. The corresponding $D_{2}$ obtained from IPR analysis (orange cross) is displayed for comparison.
@@
So far, we have considered an idealized fractal structure. In reality, random defects existing on top of the regular fractal lattice are likely unavoidable. Such imperfections can disrupt the self-similar pattern, potentially altering the nature of the NEM states observed in the pristine system. Therefore, the stability of these states against such disorder needs to be addressed.
@@
[…]
@@
According to the single-parameter scaling theory of Anderson localization [57], random geometric disorder predicts the localization of all electronic states in the thermodynamic limit for systems of dimensionality $d\leq 2$. The SG, with its fractal dimension $d_{f}\simeq 1.585$, falls into this category, implying that its eigenstates are, in principle, susceptible to localization by random perturbations. Our aim here is more specific: to assess the resilience of the previously identified NEM states against these ARDs.
@@
Figure 8 displays the $\text{nIPR}_{2}$ distribution for all eigenstates of the $g=6$ SG as a function of $N_{\mathcal{R}}$, averaged over $p=10$ geometric disorder realizations. As more ARDs are introduced, the average $\text{nIPR}_{2}$ value across the spectrum, $\langle\text{nIPR}_{2}\rangle$, exhibits a clear decreasing trend, indicating a general shift towards stronger localization. Notably, states with critical $\text{nIPR}_{2}$ values persist across certain energy intervals, even when a substantial number of ARDs are present (e.g., $N_{\mathcal{R}}=200$, corresponding to approximately $10.7\%$ of the SG sites). Once again, the underlying hierarchical structure of the SG appears to play a role in delaying a complete transition to universal Anderson localization, thereby preserving, in some states, the multifractal signature.
@@
Figure 8: Energy spectrum of the single-particle tight-binding model on the sixth-generation of Sierpiński Gasket with additional random defects (ARDs), averaged over 10 disorder realizations. Configurations with 50, 100, 150 and 200 ARDs are analysed. The color and marker size denote $\text{nIPR}_{2}$ similarly to Fig. 2. In the lower right corner of each subplot, the maximum of $\text{nIPR}_{2}$ is displayed, quantifying the tendency to localization as the number of defects increase.
@@
The overall tendency of critical states to endure upon introducing random defects can be complemented by a more detailed analysis of the robustness of certain NEMs at specific energies. While disorder-induced spectral phenomena can facilitate the reorganization of NEMs through resonances and lead new forms of multifractality, other NEMs maintain their spatial distribution without significant alteration.
@@
A remarkable example is the eigenstate at $E=1.000$ depicted in Fig. 9a, for the $6$-th generation SG. Characterized as multifractal (details in SM Sec. E), its unique wavefunction intensity profile is notably sparse, confined to several disconnected groups of sites. These clusters are predominantly located near the internal edges of the SG structure, between the third and fourth hierarchical levels. More importantly, this distinct spatial profile is reminiscent of compact localized states [58] of topologically equivalent nodes from the graph theory [59], suggesting these type of states can hold considerable potential for applications in quantum information storage and transfer [60, 61].
@@
These isolated wavefunction clusters, indeed, remain largely unaffected by ARDs unless the defects are placed directly onto their supporting lattice sites. When defects are systematically introduced on sites within these specific clusters, the degeneracy of the state is lifted, and the wavefunction amplitude is redistributed among the remaining, unaffected sets, with minimal leakage to other non-resonant sites in the lattice (Fig. 9b-d). This targeted response demonstrates a pathway for controlled manipulation of the wavefunction’s intensity, offering avenues for precise quantum control and motivating exploration of related phenomena, such as quantum scars. Moreover, as elaborated in the SM, even when a larger number of ARDs significantly perturbs the original profile, these disturbances can trigger new quantum interference effects. Consequently, amplitude from a disrupted set can re-cohere and stabilize around lower hierarchical levels of the fractal, giving rise to novel multifractal clustered patterns. The prospect of experimentally realizing these structurally unique and potentially applicable states—for instance, in atomic quantum simulators [62, 49] through precise chemical potential tuning—makes future inquiry particularly compelling.
@@
Figure 9: Amplitude profile of the state at energy level $E=1.000$. Profile (a) in the original model without defects, and (b-d) with additional defects (marked by pink crosses) placed at three, six, and eight targeted sites, respectively. Each defect, positioned within its respective cluster, completely blocks information propagation in that region, while causing lossless redistribution of information among the remaining clusters. The colorbar indicates of $|\psi|^{2}$, with red corresponding to high amplitude values and black to nearly zero ones.
@@ Discussion
With this work, we continue the research line of studying quantum phenomena in artificial fractal geometries, focusing on the case of the Sierpiński Gasket. We have shown that in clean fractals, without any conventional randomness, non-ergodic multifractal states naturally emerge, complementing prior studies about wavefunction localization [34], quantum transport [63], eigenstate statistics [40], and energy spectra [64] in fractals. We performed a multi-tool analysis of both the overall properties of the SG spectrum and the specific features of individual NEMs and provided sensitive indicators to distinguish among their diverse attributes.
@@
Among other findings, we demonstrated a particular class of NEM states (e.g., at $E=1.000$) which, to our understanding, is unique to fractal geometries. These states respond in a highly symmetric and structured way to the introduction of geometrical disorder, making this responsiveness an interesting platform for quantum control. We speculate that the displayed properties of such NEMs are analogous to those of compact localized states [58, 60] and that dispersionless wave propagation may be accomplished. This characteristic, if confirmed, would be instrumental in achieving controlled information flow between targeted regions of the system, offering new approaches to quantum control.
@@
Not to limit ourselves to the presented model, we generalized our findings (SM, Sec. G) beyond the basic SG model with only nearest-neighbor hopping. In doing so, we included additional interactions, such as next-nearest-neighbor (NNN) hopping terms and revealed that NEMs still emerge under these conditions and exhibit robustness even as the NNN hopping strength is gradually increased.
@@
Our exploration also covered other fractal structures, notably the negative Vicsek-saltire fractal (SM, Sec. H.1). In this geometry—where vacancies, rather than the lattice sites themselves, form a fractal pattern with a relatively low Hausdorff dimension ($d_{f}\simeq 1.46$)—next-nearest-neighbor hopping was also included. The main reason for considering this particular example was that its creation demands the introduction of a minimal number of defects compared with structures like the SG or the Sierpiński Carpet, making its experimental realization easier. While the large translationally invariant "clean" sub-domains support extended states, critical states still emerge in certain parts of the spectrum.
@@
We suggest that artificial fractal structures – systems that have already been created in a lab – are a natural platform to create and study multifractal states experimentally, which so far – to our knowledge – has been possible only in doped semiconductors [65] and still remains an ongoing challenge otherwise. Our numerical results, particularly on the two-point density-density correlation function $\mathcal{F}(\omega,E)$, provide a direct method to distinguish NEMs using scanning probe microscopy techniques (recent research has also demonstrated applicability of similar methods in acoustic metamaterials [66] and photonic systems [33, 67]).
@@
The signatures of multifractal states that we have analyzed here, lay the natural ground for extending this research. In particular, moving beyond the single-particle framework, future studies could reveal connections to timely many-body phenomena, such as many-body localization [68], Hilbert space fragmentation [22], and quantum scars [69, 70].

### v4

@@ Results
Left panel: Singularity spectrum $f(\alpha)$ for the representative states: extended (GS, red), localized (LOC, blue), and multifractal (NEM, orange). The spectrum for the Gaussian Orthogonal Ensemble (GOE, black line) is displayed as a reference for ideal extended states. Differences in the peak position and the width of the $f(\alpha)$ curve distinguish the localization classes. Right panel: Finite-size scaling of $f(\alpha)$ for the NEM case study. Data comparing different generation ranges show agreement, supporting its multifractal nature.
@@
Through the preceding finite-size scaling analysis of these exemplary cases, one can examine the localization properties across different groups of eigenstates, potentially unveiling a variety of degrees of multifractality. The SM Sec. D also contains data for other NEMs, displaying different types of multifractal behavior, evident in their $f(\alpha)$ spectra—for example, in the width or the shift of the peak. Some of these NEMs retain metallic or insulating characteristics [55, 56] and may also exhibit spatial structures that resemble bulk and edge modes observed in artificial fractal systems [34, 35].
@@
It is worth emphasizing that, unlike Random Matrix Theory systems, which require ensemble averaging over disorder realizations, our system is deterministic and inherently disorder-free. Thus, $\mathcal{F}(\omega,E)$ directly measures the intrinsic correlations for our specific Hamiltonian between the densities of eigenstates $\psi_{i}$ and $\psi_{j}$ (with eigenvalues $\lambda_{i}$ and $\lambda_{j}$, respectively) that are separated by an energy $\omega$, with their average energy $(\lambda_{i}+\lambda_{j})/2$ centered around $E$. This approach provides an experimentally accessible probe to reveal and discriminate between extended, localized, and NEM signature states, particularly through its characteristic scaling behavior at criticality.
@@
Figure 6 illustrates $\mathcal{F}(\omega,E)$ for the previously selected representative extended, localized, and NEM states, computed for the $6$th generation SG. As predicted by theory, the two-point correlator for the extended state (red) and the localized state (blue) exhibits a plateau across several orders of magnitude at small energy separation in $\omega$. The plateau height reflects the degree of eigenstate overlap: relatively low and broad for extended states, due to their diffuse nature across the system, and significantly higher for localized states, which are concentrated in confined regions. At large energy separations in $\omega$, the correlator approaches the uncorrelated regime, with exponential decay for localized states. In this large-$\omega$ limit, $\mathcal{F}(\omega,E)$ tends towards unity (i.e., $\mathcal{F}\approx 1$), a value consistent with representative GOE statistics.
@@
Figure 6: Log-log plot of the two-point density-density correlator $\mathcal{F}(\omega,E)$ versus energy separation $\omega$. Data are shown for three characteristic states from a 66th generation Sierpiński gasket: the extended ground state (GS, $E\simeq-3.667$, red markers), a localized state (LOC, $E\simeq-2.184$, blue markers), and a non-ergodic multifractal state (NEM, $E\simeq 1.509$, orange markers). The dashed black line indicates the Gaussian Orthogonal Ensemble (GOE) reference for uncorrelated systems, corresponding to $\mathcal{F}(\omega,E)\approx 1$. While both extended and localized states exhibit broad plateaus at small $\omega$ indicating non-vanishing correlations, their respective heights are markedly different. Specifically, for the chosen localized state—situated in a spectral gap—the high plateau is a direct consequence of its triple degeneracy. In contrast, the NEM state features a narrower one (expected to become a singular point in the thermodynamic limit) followed by a characteristic power-law decay, $\mathcal{F}(\omega,E)\propto\omega^{-\mu}$, a fingerprint of multifractality. Lines are guides to the eye.
@@
The two-point correlator for the NEM state (orange) displays a distinctly shorter plateau, followed by a characteristic power-law decay with energy separation $\omega$, described by $\mathcal{F}(\omega,E)\propto\omega^{-\mu}$, before approaching the uncorrelated regime. Power-law scaling over a broad $\omega$ range is a hallmark of multifractal states and reflects their inherent scale-invariance [55, 58].
@@
To quantitatively verify this multifractal scaling and determine the correlation dimension $D_{2}$, we analyze the system-size dependence of $\mathcal{F}(\omega,E;N)$. The left panel of Figure 7 presents data for the NEM state at $E\simeq 1.509$ across multiple system sizes. As the system size $N$ increases, the $\mathcal{F}(\omega,E;N)$ curves exhibit improved convergence in their power-law region.
@@
Based on Chalker’s scaling hypothesis for critical systems [55], the exponent $\mu(N)$ is related to a size-dependent correlation dimension $D_{2}(N)$ by $\mu(N)\approx 1-D_{2}(N)$. Thus, the value for $D_{2}$ in the thermodynamic limit is obtained by extrapolating these finite-size $D_{2}(N)$ values against $1/\ln N$. The finite-size scaling data are plotted in the right panel of Fig. 7 and yield the fractal dimension $D_{2}=0.661\pm 0.012$. This value is in agreement, within the stated uncertainties, with the value $D_{2}\simeq 0.652$ obtained independently from IPR-based eigenstate statistics. This consistency between two distinct methods strongly corroborates the emergence of multifractal states.
@@
Left panel: Log-log plot of $\mathcal{F}(\omega,E;N)$ versus energy separation $\omega$ for Sierpiński gasket generations g=57g=5-7 (circles, solid lines, exact diagonalization) and g=89g=8-9 (triangles, dashed lines, subspace diagonalization). The dotted black line shows the extrapolated power-law decay (with exponent $\mu$ from the thermodynamic-limit value of $D_{2}$ in the right panel), while the gray shaded area marks the regime used for the fits. The dashed black line represents the GOE uncorrelated limit ($\mathcal{F}=1$). Right panel: Finite-size scaling extrapolation of the correlation dimension $D_{2}(N)\approx 1-\mu(N)$ against $1/\ln(N)$. Blue circles denote the $D_{2}(N)$ values derived from the decay exponents $\mu(N)$ for each generation. The dash-dot line is a linear fit, with its $y$-intercept (red circle) collapsing to thermodynamic limit value $D_{2}=0.661\pm 0.012$. Statistical errors are included. The corresponding $D_{2}$ obtained from IPR analysis (orange cross) is displayed for comparison.
@@
So far, we have considered an idealized fractal structure. In reality, random defects existing on top of the regular fractal lattice are likely unavoidable. Such imperfections can disrupt the self-similar pattern, altering the nature of the NEM states observed in the pristine system. Therefore, the stability of these states against such disorder needs to be addressed.
@@
$H_{\text{ARD}}=\!\!\sum\limits_{\begin{subarray}{c}\langle i,j\rangle\\ i,j\notin(\mathcal{B}\cup\mathcal{R})\end{subarray}}\!\!t\,c^{\dagger}_{i}c_{j}\,\,+\!\sum_{k\in(\mathcal{B}\cup\mathcal{R})}\!\!V\,c^{\dagger}_{k}c_{k}\,\,+\!\!\sum_{\begin{subarray}{c}\langle i,k\rangle\\ k\in(\mathcal{B}\cup\mathcal{R})\end{subarray}}\!\!\overline{t}\,c^{\dagger}_{i}c_{k}+\text{h.c.}$ (8)
@@
According to the single-parameter scaling theory of Anderson localization [59], random geometric disorder leads to the localization of all electronic states for systems of dimensionality $d\leq 2$. The SG, with its fractal dimension $d_{f}\simeq 1.585$, falls into this category, implying that its eigenstates would ultimately localize under such perturbations in the thermodynamic limit. Our aim here is more specific: to assess how the previously identified NEM states are affected by ARDs, and to what extent their characteristic properties survive in experimentally relevant finite-size systems.
@@
Figure 8 displays the $\text{nPR}_{2}$ distribution for all eigenstates of the $g=6$ SG as a function of $N_{\mathcal{R}}$, averaged over $p=10$ geometric disorder realizations. As more ARDs are introduced, the average $\text{nPR}_{2}$ value across the spectrum, $\langle\text{nPR}_{2}\rangle$, exhibits a clear decreasing trend, indicating a general shift towards stronger localization. Notably, critical states remain present across certain energy intervals, even when a substantial number of ARDs are present (e.g., $N_{\mathcal{R}}=200$, corresponding to approximately $10.7\%$ of the SG sites). Once again, the underlying hierarchical structure of the SG seems to delay a complete transition to universal Anderson localization, thereby preserving, in some states, the multifractal signature.
@@
Figure 8: Energy spectrum of the single-particle tight-binding model on the 6th generation of Sierpiński gasket with additional random defects (ARDs), averaged over 10 disorder realizations. Configurations with 50, 100, 150 and 200 ARDs are analyzed. The color and marker size denote $\text{nPR}_{2}$ similarly to Fig. 2. In the lower right corner of each subplot, the maximum of $\text{nPR}_{2}$ is displayed, quantifying the tendency to localization as the number of defects increases.
@@
The overall tendency of critical states to endure upon introducing random defects can be complemented by a more detailed analysis of the robustness of certain NEMs at specific energies. While disorder-induced spectral phenomena can facilitate the reorganization of NEMs through resonances and lead to new forms of multifractality, other NEMs maintain their spatial distribution without significant alteration.
@@
A remarkable example is the eigenstate at $E\simeq 1.000$ depicted in Fig. 9a, for the $6$th generation SG. Characterized as multifractal (details in SM Sec. E), its unique wavefunction intensity profile is sparse, confined to several disconnected groups of sites. These clusters are predominantly located near the internal edges of the SG structure, between the third and fourth hierarchical levels. More importantly, this distinct spatial profile is reminiscent of compact localized states [60] and of topologically equivalent nodes from the graph theory [61], suggesting these types of states can hold considerable potential for applications in quantum information storage and transfer [62, 63].
@@
These isolated wavefunction clusters, indeed, remain largely unaffected by ARDs unless the defects are placed directly onto their supporting lattice sites. When defects are systematically introduced on sites within these specific clusters, the degeneracy is lifted, and the wavefunction amplitude is redistributed among the remaining, unaffected sets, with minimal leakage to other non-resonant sites (Fig. 9b-d). This targeted response demonstrates a pathway for controlled manipulation of the wavefunction’s intensity, offering avenues for precise quantum control and motivating exploration of related phenomena, such as quantum scars. Moreover, as elaborated in the SM Sec. F, even when a larger number of ARDs significantly perturbs the original profile, these disturbances can trigger new quantum interference effects. Consequently, amplitude from a disrupted set can re-cohere and stabilize around lower hierarchical levels of the fractal, giving rise to novel multifractal clustered patterns. The prospect of experimentally realizing these structurally unique and potentially applicable states—for instance, in atomic quantum simulators [64, 49] through precise chemical potential tuning—makes future inquiry particularly compelling.
@@
Figure 9: Amplitude profile of the state at energy level $E=1.000$. Profiles (a) in the original model without defects, and (b-d) with additional defects (marked by pink crosses) placed at three, six, and eight targeted sites, respectively. Each defect, positioned within its respective cluster, completely blocks information propagation in that region, while causing lossless redistribution of information among the remaining clusters. The colorbar indicates $|\psi|^{2}$, with red corresponding to high amplitude values and black to nearly zero ones.
@@ Discussion
With this work, we continue the study of quantum phenomena in artificial fractal geometries, focusing on the Sierpiński gasket. We have shown that in clean fractals, without any conventional randomness, non-ergodic multifractal states naturally emerge, complementing prior studies on wavefunction localization [34], quantum transport [65], eigenstate statistics [40], and energy spectra [66] in fractals. We performed a multi-tool analysis of the properties of the SG spectrum and the features of individual NEMs. In addition, we provided sensitive indicators that distinguish their diverse attributes, confirming that these states are robust and generic rather than rare (SM, Sec. F).
@@
Among other findings, we demonstrated a particular class of NEM states (e.g., at $E\simeq 1.000$) which, to our understanding, is unique to fractal geometries. These states respond in a highly symmetric and structured way to the introduction of geometric disorder, making this responsiveness an interesting platform for quantum control. We speculate that the displayed properties of such NEMs are analogous to those of compact localized states [60, 62] and that dispersionless wave propagation may be accomplished. If confirmed, this could enable controlled information flow between targeted regions, offering new approaches to quantum control.
@@
Not to limit ourselves to the presented model, we generalized our findings beyond the basic SG model with only nearest-neighbor hopping. In doing so, we included additional interactions, such as next-nearest-neighbor (NNN) hopping terms and revealed that NEMs still emerge under these conditions and exhibit robustness even as the NNN hopping strength is gradually increased (SM, Sec. F.3).
@@
Our exploration also covered other fractal structures, notably the negative Vicsek-saltire fractal (SM, Sec. G.1). In this geometry, vacancies, rather than the lattice sites themselves, form a fractal pattern with a relatively low Hausdorff dimension ($d_{f}\simeq 1.46$). This particular example was chosen because its creation demands fewer defects than the SG or the Sierpiński carpet, making its experimental realization easier. While the large translationally invariant "clean" sub-domains support extended states, critical states still emerge in certain parts of the spectrum.
@@
We suggest that artificial fractal structures—systems that have already been created in a lab—are a natural platform to create and study multifractal states experimentally, which, to our knowledge, has so far been possible only in doped semiconductors [67] and still remains an ongoing challenge. Our numerical results, particularly on the two-point density-density correlation function $\mathcal{F}(\omega,E)$, provide a direct method to distinguish NEMs using scanning probe microscopy techniques (recent research has also demonstrated applicability of similar methods in acoustic metamaterials [68] and photonic systems [33, 69]).
@@
The signatures of multifractal states that we have analyzed here lay the natural ground for extending this research. In particular, moving beyond the single-particle framework, future studies could reveal connections to timely many-body phenomena, such as many-body localization [70], Hilbert space fragmentation [22], and quantum scars [71, 72].